Jill asks one of her classmates to choose a number, then multiply the number by -3, add 2 to the product, multiply the result by -2, and then subtract 14. Finally, Jill asks her classmate to divide the result by 6 and record the answer.
When Jill gets the answer from her classmate, she just adds 3 to it in her head and announces the number that her classmate originally chose. How did Jill know to add 3 to her classmate's answer?

Answer by htmentor(1343)

Let x = the number chosen
The operations that get applied to the number are:
(-2(-3*x + 2) - 14 [-3 times the number, add 2, multiply everything by -2, then subtract 14]
This gives 6x - 4 - 14 = 6x - 18
Now divide by 6:
(6x-18)/6 = x - 3
If we now add 3 to this expression we get the original number, x